# Making Machines Sound Sarcastic: LLM-Enhanced and Retrieval-Guided Sarcastic Speech Synthesis

### v1

@@ Abstract
Sarcasm is a subtle form of non-literal language that poses significant challenges for speech synthesis due to its reliance on nuanced semantic, contextual, and prosodic cues. While existing speech synthesis research has focused primarily on broad emotional categories, sarcasm remains largely unexplored. In this paper, we propose a Large Language Model (LLM)-enhanced Retrieval-Augmented framework for sarcasm-aware speech synthesis. Our approach combines (1) semantic embeddings from a LoRA-fine-tuned LLaMA 3, which capture pragmatic incongruity and discourse-level cues of sarcasm, and (2) prosodic exemplars retrieved via a Retrieval Augmented Generation (RAG) module, which provide expressive reference patterns of sarcastic delivery. Integrated within a VITS backbone, this dual conditioning enables more natural and contextually appropriate sarcastic speech. Experiments demonstrate that our method outperforms baselines in both objective measures and subjective evaluations, yielding improvements in speech naturalness, sarcastic expressivity, and downstream sarcasm detection.
Index Terms—  Sarcastic speech synthesis, sarcasm detection, large language models, retrieval augmented generation
1 Introduction
Sarcasm is a common yet challenging form of non-literal language, often marked by prosodic patterns such as exaggerated intonation and prolonged syllables [1, 2]. Its proper interpretation relies on a nuanced interplay of semantic, contextual, and prosodic cues, making it difficult to capture computationally. While sarcasm is widely used in everyday communication, it has received limited attention in the field of speech synthesis. Most existing work on expressive text-to-speech (TTS) has concentrated on broad emotional categories such as happiness, anger, or sadness [3, 4, 5], whereas more subtle and pragmatically grounded attitudes such as sarcasm remain largely unexplored. While prior work has attempted to analyze acoustic correlates of sarcastic speech, only a limited number of studies have explored directly manipulating prosodic features such as pitch, pace, and loudness to generate sarcastic-sounding speech [6]. In addition, current TTS systems, while successful at producing intelligible and emotionally expressive speech, often lack the nuanced expressiveness needed to convey sarcastic intent [5, 7, 8].
Sarcastic speech synthesis holds considerable potential for improving human-computer interaction in applications such as conversational agents and entertainment systems [9]. However, the task presents two major challenges. First, high-quality, annotated sarcastic speech corpora remain scarce, limiting the applicability of purely data-driven approaches. Existing corpora for sarcasm, such as MUStARD [10] or its extensions [11], are primarily developed for detection rather than synthesis. Second, sarcasm is inherently more subtle and context-dependent than conventional emotions such as anger or joy, making it difficult to capture with handcrafted acoustic features or simple prosodic controls. In contrast, research on sarcasm detection has progressed significantly. Recent work has shown that multi-modal approaches, incorporating textual, prosodic and visual information, can significantly improve sarcasm recognition [10, 11, 12, 13, 14]. These advances highlight the feasibility of modeling sarcasm computationally. Yet, most efforts have remained on the recognition side, leaving the generation of sarcastic speech largely unaddressed. Moving from detection to synthesis is not only a natural next step but also a crucial one, as it enables interactive systems to actively deploy pragmatic phenomena rather than passively identify them.
@@ 1 Introduction
To address these limitations, recent advances in natural language processing offer promising directions. LLMs have demonstrated the ability to capture high-level semantic and pragmatic features from text. In the TTS domain, integrating LLM-derived embeddings has been shown to enable more nuanced prosody control and improve emotional expressiveness and contextual appropriateness [15, 16, 17]. Such capabilities are particularly relevant for sarcasm, which often relies on pragmatic incongruity between literal meaning and intended attitude. Leveraging LLM-based semantic representations could therefore provide a principled way of conditioning TTS models on sarcasm-relevant cues that handcrafted features fail to capture.
@@
Complementary to this, RAG allows models to condition generation on external knowledge or examples retrieved from large databases [18, 19]. While RAG has primarily been explored in text-based tasks such as question answering and dialogue [20, 21], its potential for speech synthesis is evidenced by related work on style transfer and expressive TTS, where reference utterances are used to guide prosody [18]. Existing approaches, however, typically rely on manually selected or speaker-specific reference audio, limiting scalability and contextual alignment. In the case of sarcasm, where annotated corpora are scarce, automatic retrieval of semantically similar sarcastic speech could provide prosodic exemplars that both enrich training and guide generation. This suggests a paradigm in which LLMs supply semantic-pragmatic cues while RAG retrieves expressive references, together enabling a systematic modeling of subtle pragmatic phenomena such as sarcasm in speech synthesis.
Fig. 1: Overview of the sarcasm-aware speech synthesis framework. Semantic cues from a LoRA fine-tuned LLaMA 3 and prosodic exemplars retrieved via RAG are fused within a VITS-based model to synthesize sarcastic speech.
This study introduces a unified sarcasm-aware speech synthesis framework that enriches the TTS system with LLM-extracted semantic cues indicative of sarcasm and prosodically aligned audio references for expressive modeling. Specifically, we first use an LLM to extract high-level semantic representations capturing features relevant to sarcastic intent. In parallel, a retrieval mechanism based on RAG identifies semantically similar and contextually aligned prosodic examples from a curated speech database, providing prosodic exemplars that guide the TTS model toward expressive and contextually appropriate output. Both objective metrics and subjective evaluations demonstrate that the proposed method outperforms baseline models in naturalness, expressivity, and accurate sarcasm expression.
2 Methods
@@
To achieve high-quality sarcastic speech synthesis, we propose a Retrieval-Augmented LLM-enhanced TTS framework. An overview of the architecture is illustrated in Figure 1.
@@
2.1 Sarcasm-Aware Semantic Encoding via LoRA
@@
Low-Rank Adaptation (LoRA) has recently been widely adopted for adapting LLMs across diverse domains such as emotion recognition, sentiment analysis, and affective detection, due to its parameter efficiency and effectiveness [22]. Motivated by these advances, we leverage LoRA to finetune the LLaMA 3-8B model for sarcasm-aware semantic encoding. Fine-tuning is performed on a curated sarcasm-labeled dataset, enabling the model to capture signals of sarcastic intent, including pragmatic incongruity and discourse-level cues.
@@
Formally, given an input text $x$, the Sarcasm-LoRA encoder produces a semantic embedding:
@@
where $d_{t}$ denotes the dimensionality of the semantic embedding. These embeddings capture both contextual and pragmatic aspects of the input, providing a sarcasm-aware representation for further processing.
2.2 RAG for Prosody Conditioning
To enhance prosodic expressiveness, we integrate a RAG module into the synthesis pipeline. The key idea is to leverage semantically aligned sarcastic utterances as prosodic exemplars, enriching the generated speech with realistic intonation patterns. Concretely, we first construct an index $\mathcal{D}=\{(u_{i},\mathbf{a}_{i})\}_{i=1}^{N}$ of sarcastic utterances $u_{i}$ curated from MUStARD++, where $\mathbf{a}_{i}$ denotes their pre-computed semantic representations. For a given input text, the semantic embedding $\mathbf{E}_{s}$ produced by the Sarcasm-LoRA encoder is used as a retrieval query to fetch the top-$K$ semantically and pragmatically relevant sarcastic utterances from the indexed database.
@@ 2.2 RAG for Prosody Conditioning
where $\text{Pool}(\cdot)$ aggregates the frame-level features into a fixed-length vector. These exemplars provide style references that reflect the characteristic intonation and emphasis patterns of sarcastic speech.
Compared to conventional style transfer methods that rely on a single manually chosen reference utterance, the proposed retrieval mechanism offers two key advantages. First, it scales more naturally to diverse input contexts, since the retrieval process automatically identifies semantically aligned exemplars. Second, by conditioning on multiple retrieved samples, the model can capture a richer variety of sarcastic prosodic patterns, leading to more robust and contextually appropriate expressive synthesis.
2.3 LLM-Enhanced and Retrieval-Guided Sarcastic TTS
@@
Let $\mathbf{E}_{p}\in\mathbb{R}^{T_{p}\times d_{p}}$ denote the phoneme embeddings and $\mathbf{E}_{s}\in\mathbb{R}^{T_{t}\times d_{t}}$ the sarcasm-aware semantic embeddings produced by the fine-tuned LLaMA 3. In the cross-attention module, we treat $\mathbf{E}_{p}$ as the query and $\mathbf{E}_{s}$ as key and value:
@@
where $\mathbf{W}_{q},\mathbf{W}_{k},\mathbf{W}_{v}$ are learned projection matrices, and $d_{k}$ is the dimensionality of the key.
Finally, the prosody exemplars $\mathcal{E}_{w}=\{\mathbf{E}_{w_{1}},\dots,\mathbf{E}_{w_{K}}\}$ extracted by WavLM are linearly projected and added to the cross-attention output to modulate the decoder:
@@
3 Experiments
To evaluate the effectiveness of our proposed approach to sarcastic speech synthesis, we conducted a series of experiments involving baseline, LLM-enhanced, RAG-enhanced, and combined models. These experiments assess the impact of semantic guidance from language models and reference-based prosodic conditioning on the quality and sarcasm-awareness of synthesized speech [^1] Speech samples are available at https://abel1802.github.io/RAG-LLM-SarcasticTTS/.
3.1 Data
For the initial pre-training phase of VITS, we use the HiFi-TTS corpus [25], a high-fidelity audiobook-derived dataset with diverse speakers and wideband, clean recordings.
@@ 3.1 Data
To adapt LLMs as sarcasm-aware semantic encoders for speech synthesis, we fine-tune them on the News Headlines Sarcasm dataset [26]. This dataset contains over 28,000 headlines from sources such as The Onion, each annotated with a binary sarcasm label. The class distribution is roughly balanced, making it suitable for learning discourse-level markers of sarcastic intent. The LLM serves as a feature extractor that provides high-level semantic embeddings conditioned on sarcastic intent. To validate the effectiveness of this adaptation, we also evaluate the fine-tuned model on a sarcasm detection task on the whole MUStARD++ dataset [11], in order to test that the learned embeddings capture sarcasm-relevant cues.
3.2 Experimental Setup
@@
3.3 Compared Methods
We compare the proposed framework against a series of baseline and enhanced variants to systematically evaluate the contribution of different modules:
  • VITS (Baseline): A standard variational text-to-speech model serving as the baseline, without explicit semantic or stylistic guidance.
  • VITS + BERT (w/ BERT): Extends VITS with semantic embeddings from a pretrained BERT model [27], providing additional semantic features during synthesis.
  • VITS + LLaMA 3 (w/ LLaMA 3): Uses textual embeddings generated by LLaMA 3 [28], enabling richer contextual and pragmatic guidance.
  • VITS + LLaMA 3-LoRA (w/ LoRA): Incorporates a LoRA-fine-tuned version of LLaMA 3 trained on sarcasm-labeled data, enhancing the ability to capture sarcasm-relevant semantics.
  • VITS + RAG (w/ RAG): Employs RAG to retrieve sarcastic reference speech samples that are semantically similar to the input text. These references are used to provide prosodic guidance during synthesis.
  • VITS + LLaMA 3-LoRA + RAG (Proposed): Our proposed unified framework, combining sarcasm-aware semantic embeddings from fine-tuned LLaMA 3 with prosodic exemplars retrieved through RAG. This model is designed to maximize both semantic relevance and expressive alignment with sarcastic intent.
4 Results and Discussion
4.1 Sarcasm Detection Results
We evaluate sarcasm detection performance from text-only inputs on the MUStARD++ dataset, comparing the original sarcasm detection systems used in MUStARD++ [11], BERT, LLaMA 3, and LLaMA 3-LoRA. Models are evaluated with Precision (P), Recall (R), and weighted F1-score (F1), with results summarized in Table 1.
Table 1: Performance of different models on sarcasm detection (text-only input).
@@ 4.1 Sarcasm Detection Results
Table 2: Evaluation results of TTS systems. MCD is in dB; pitch and energy are averaged with standard deviation; Sarcasm detection includes precision (P), recall (R), and weighted F1-score (F1); MOS and SMOS denote naturalness and sarcasm MOS scores, respectively.
Objective Sarcasm Detection Subjective
Method MCD $\downarrow$ Pitch $\downarrow$ Energy $\downarrow$ P (%) $\uparrow$ R (%) $\uparrow$ F1 (%) $\uparrow$ MOS $\uparrow$ SMOS $\uparrow$
GT 62.8 62.3 62.3 $3.8\pm 0.1$ $4.5\pm 0.2$
Baseline $9.8\pm 3.0$ $261.9\pm 148.8$ $4.4\pm 1.4$ 60.3 60.5 59.9 $2.6\pm 0.1$ $3.2\pm 0.2$
w/ BERT $\mathbf{9.6\pm 2.8}$ $265.6\pm 147.1$ $4.5\pm 1.4$ 60.5 60.6 60.5 $2.5\pm 0.1$ $3.1\pm 0.2$
w/ LLaMA 3 $10.4\pm 3.1$ $261.6\pm 135.1$ $4.4\pm 1.0$ 59.6 59.7 59.0 $2.0\pm 0.1$ $2.6\pm 0.2$
w/ LoRA $10.1\pm 3.1$ $282.6\pm 149.6$ $4.4\pm 1.4$ 60.6 60.8 60.6 $2.6\pm 0.1$ $3.8\pm 0.2$
w/ RAG $10.0\pm 3.0$ $261.0\pm 148.0$ $4.4\pm 1.4$ 61.5 61.7 61.6 $2.6\pm 0.1$ $3.7\pm 0.2$
Proposed $9.8\pm 2.8$ $\mathbf{259.6\pm 152.3}$ $\mathbf{4.4\pm 1.4}$ 62.7 62.9 62.5 $\mathbf{2.7\pm 0.1}$ $\mathbf{3.8\pm 0.2}$
@@
As shown in Table 1, both the MUStARD++ baseline and BERT achieved moderate performance with F1-scores around 67%. LLaMA 3 performed slightly worse (65.5%), suggesting that general-purpose pretraining alone is insufficient to capture the nuanced cues of sarcasm. In contrast, parameter-efficient fine-tuning with LoRA yielded a clear improvement, raising the F1-score to 72.5%. This demonstrates the effectiveness of the proposed sarcasm-aware semantic encoding via LoRA (Section 2.1), and motivates its integration into the sarcasm-aware TTS framework to further examine whether such improvements in semantic encoding can also lead to more expressive synthesized speech.
4.2 Sarcastic Speech Synthesis Results

### v2

# Modeling Sarcastic Speech: Semantic and Prosodic Cues in a Speech Synthesis Framework

@@ Abstract
Sarcasm is a pragmatic phenomenon in which speakers convey meanings that diverge from literal content, relying on an interaction between semantics and prosodic expression. However, how these cues jointly contribute to the recognition of sarcasm remains poorly understood. We propose a computational framework that models sarcasm as the integration of semantic interpretation and prosodic realization. Semantic cues are derived from an LLaMA 3 model fine-tuned to capture discourse-level markers of sarcastic intent, while prosodic cues are extracted through semantically aligned utterances drawn from a database of sarcastic speech, providing prosodic exemplars of sarcastic delivery. Using a speech synthesis testbed, perceptual evaluations demonstrate that both semantic and prosodic cues independently enhance listeners’ perception of sarcasm, with the strongest effects emerging when the two are combined. These findings highlight the complementary roles of semantics and prosody in pragmatic interpretation and illustrate how modeling can shed light on the mechanisms underlying sarcastic communication.
Keywords: pragmatics; sarcasm; prosody; semantic cues; speech perception
Introduction
Sarcasm plays an important role in everyday communication by enabling speakers to express mockery, criticism, or contempt through meanings that diverge from literal interpretation. As a pragmatic phenomenon, sarcasm is often realized through irony, where ostensibly positive or neutral expressions convey an underlying negative intent. While such ironic intent can sometimes be inferred from explicit linguistic content, sarcasm frequently lacks clear lexical markers and therefore cannot be reliably identified from semantics alone. Instead, its interpretation depends on a nuanced interplay of semantic incongruity, contextual information, and prosodic cues. Speakers commonly employ prosodic strategies, such as exaggerated intonation, emphatic stress, or lengthened syllables, to signal sarcastic intent, particularly in cases where linguistic cues are subtle or ambiguous [27, 7, 18]. This reliance on multiple cues makes sarcasm a complex phenomenon that challenges computational modeling, as successful automatic recognition requires integrating information across modalities. Understanding how listeners integrate these cues is a fundamental question for theories of pragmatic language comprehension.
One challenge in studying sarcastic communication lies in systematically manipulating semantic and prosodic cues while holding other factors constant. Traditional corpus-based or behavioral studies often lack fine-grained control over semantic and prosodic cues, since these dimensions tend to be entangled in natural speech and are difficult to manipulate independently or in a graded manner as experimental stimuli [3, 4, 28]. Speech synthesis, however, offers a powerful experimental testbed, enabling controlled generation of utterances in which semantic and prosodic information can be independently varied to explore their respective contributions to sarcasm perception.
Recent advances in computational modeling provide new opportunities to formally represent and manipulate the cues underlying pragmatic communication. Large language models (LLMs) have shown an ability to capture discourse-level semantic and pragmatic patterns, including forms of semantic subtlety relevant to sarcasm. At the same time, retrieval-based approaches allow models to condition generation on prosodic exemplars that reflect expressive sarcasm patterns. Together, these methods offer a way to formalize hypotheses about how semantic content and prosodic realization jointly contribute to sarcastic communication.
In this study, we introduce a sarcasm-aware speech synthesis framework that models sarcasm as the integration of semantic content and prosodic expression. We independently manipulate two types of cues to condition sarcastic speech generation in our synthesis framework: 1) semantic representations of sarcastic intent derived from a fine-tuned LLM, and 2) prosodic patterns drawn from aligned reference utterances in a curated speech database of sarcastic speech. Using perceptual evaluations of the synthesized speech, we assess how each cue contributes to listeners’ perception of sarcasm. Perceptual evaluations reveal that both cues independently enhance sarcasm recognition, with the strongest effects emerging when they are combined. These findings provide computational evidence for the complementary roles of semantics and prosody in pragmatic interpretation and demonstrate how synthesis-based models can serve as tools for modeling speech behavior.
Related Work
Sarcasm Detection and Modeling
In recent years, substantial progress has been made in computational sarcasm detection. Many studies have adopted multimodal approaches that integrate textual, acoustic, and visual cues, demonstrating significant improvements over unimodal methods [6, 25, 24, 12, 19]. Datasets such as MUStARD [6] and its extensions [25] have facilitated this line of work by providing aligned textual, auditory, and visual data for sarcasm recognition. While these efforts highlight the feasibility of modeling sarcasm computationally, they have largely focused on classification and recognition tasks. As a result, relatively little attention has been paid to the generation of sarcastic speech or to using computational models as tools for exploring how different cues contribute to sarcasm perception. Moving from detection to synthesis is not only a natural next step but also a crucial one, as it enables interactive systems to actively deploy pragmatic phenomena rather than passively identify them.
Expressive Speech Synthesis and Sarcasm
Research on expressive text-to-speech (TTS) synthesis has primarily concentrated on broad emotional categories such as happiness, anger, and sadness [31, 1, 16]. Although these systems can produce emotionally expressive and intelligible speech, more subtle and pragmatic phenomena remain largely unexplored. Sarcasm, despite being widely used in everyday communication, has received particularly little attention in speech synthesis research. However, sarcastic or other pragmatic speech synthesis holds considerable potential for improving human-computer interaction in applications such as conversational agents and entertainment systems [26]. Prior work has attempted to analyze acoustic correlates of sarcastic speech, identifying prosodic features related to pitch, pace, and loudness [7, 18]. However, only a limited number of studies have explored directly manipulating such prosodic features to generate sarcastic-sounding speech [23]. This difficulty arises in part because sarcasm is inherently more subtle and context-dependent than conventional emotions such as anger or joy, making it difficult to capture with handcrafted acoustic features or simple prosodic controls. Moreover, the development of sarcastic speech synthesis systems is hindered by the scarcity of high-quality, annotated sarcastic speech corpora, which limits the applicability of purely data-driven approaches [17, 20].
LLM-Based Semantics and Retrieval-Based Speech Synthesis
@@ Introduction
Recent advances in natural language processing have introduced new opportunities for modeling pragmatic phenomena. LLMs have demonstrated the ability to capture high-level semantic and discourse-level information. In the TTS domain, integrating LLM-derived embeddings has been shown to enable more nuanced prosody control and improve emotional expressiveness and contextual appropriateness [11, 15, 29]. Such semantic representations are particularly relevant for sarcasm, which often depends on pragmatic incongruity between literal meaning and intended attitude. Leveraging LLM-based semantic representations could therefore provide a principled way of conditioning TTS models on sarcasm-relevant cues that handcrafted features fail to capture.
@@
Complementary to semantic modeling, retrieval-based methods have been explored as a way to condition speech generation on external knowledge or examples retrieved from large databases. Retrieval-augmented generation (RAG) has been widely applied in text-based tasks such as question answering and dialogue [14, 30], and related ideas have been adopted in speech synthesis through reference-based and style-transfer approaches, where reference utterances are used to guide prosodic realization [33, 21]. However, existing approaches often rely on manually selected or speaker-specific reference audio, limiting scalability and contextual alignment. In the context of sarcasm, where annotated corpora are scarce, automatic retrieval of semantically similar sarcastic speech could provide prosodic exemplars that both enrich training and guide generation in data-scarce settings. This suggests a paradigm in which LLMs supply semantic-pragmatic cues while the RAG module retrieves expressive references for prosodic encoding, together enabling a systematic modeling of subtle pragmatic phenomena such as sarcasm.
Figure 1: Overview of the proposed framework modeling sarcasm as an interaction between semantic and prosodic cues. Semantic representations capturing semantic and pragmatic cues are extracted from a sarcasm-adapted LLaMA 3 model, while prosodic exemplars are retrieved from reference sarcastic speech. These cues are integrated within a speech synthesis architecture to generate utterances that convey sarcastic intent.
Methods
@@
We propose a Retrieval-Augmented LLM-enhanced sarcastic speech modeling framework. An overview of the architecture is illustrated in Figure 1.
@@
Sarcasm-Aware Semantic Encoding via Parameter-Efficient Fine-tuning (PEFT)
To systematically manipulate semantic cues of sarcasm, we encode input text into sarcasm-aware semantic representations using a PEFT method on a sarcasm-labeled corpus. This allows us to isolate and represent high-level semantic features indicative of sarcastic intent for subsequent synthesis and perceptual evaluation.
@@
Low-Rank Adaptation (LoRA) has recently been widely adopted for adapting LLMs across diverse domains such as emotion recognition, sentiment analysis, and affective detection, due to its parameter efficiency and effectiveness [5]. Motivated by these advances, we leverage LoRA to finetune the LLaMA 3-8B model for sarcasm-aware semantic encoding. Fine-tuning is performed on a curated sarcasm-labeled dataset, enabling the model to capture signals of sarcastic intent, including pragmatic and discourse-level cues.
@@
Formally, given an input text sequence $x=(w_{1},\dots,w_{T_{t}})$ of length $T_{t}$, the LLaMA 3-LoRA encoder produces a sequence of contextualized semantic embeddings:
@@
where $d_{t}$ denotes the hidden dimensionality of the language model. Each row of $\mathbf{E}_{s}$ corresponds to a token-level representation extracted from the hidden state of the final transformer layer, encoding meanings relevant to sarcasm and providing a controllable semantic representation for sarcasm synthesis.
RAG for Prosody Conditioning
To investigate the role of prosodic cues in sarcasm perception, we retrieve semantically aligned sarcastic utterances as prosodic exemplars. These exemplars allow us to manipulate prosody independently of semantic content for synthesis.
We integrate an RAG module into the sarcastic speech synthesis framework. The key idea is to leverage semantically aligned sarcastic utterances as prosodic references, providing the generated speech with realistic intonation patterns. Concretely, we first construct an index $\mathcal{D}=\{(u_{i},\mathbf{a}_{i})\}_{i=1}^{N}$ of sarcastic utterances $u_{i}$ curated from MUStARD++, where $\mathbf{a}_{i}$ denotes their pre-computed semantic representations using the same LLaMA 3-LoRA adapted LLaMA 3-8B encoder. For a given input text, the semantic embedding $\mathbf{E}_{s}$ produced by the LLaMA 3-LoRA encoder is used as a retrieval query to fetch the top-$K$ semantically relevant sarcastic utterances from the indexed database; in this work, we set $K=3$.
@@
where $\text{Pool}(\cdot)$ aggregates the frame-level features into a fixed-length vector, and $d_{w}$ denotes the dimensionality of the resulting prosodic embedding. These exemplars provide style references that reflect the characteristic intonation and emphasis patterns of sarcastic speech.
Compared to conventional style transfer methods that rely on a single manually chosen reference utterance, the proposed retrieval mechanism offers two key advantages. First, it scales more naturally to diverse input contexts, since the retrieval process automatically identifies semantically aligned exemplars. Second, by conditioning on multiple retrieved samples, the model can capture a richer variety of sarcastic prosodic patterns, leading to more contextually appropriate speech.
LLM-Enhanced and Retrieval-Guided TTS for Sarcasm
@@
Let $\mathbf{E}_{p}\in\mathbb{R}^{T_{p}\times d_{p}}$ denote the sequence of phoneme embeddings, where $T_{p}$ is the number of phonemes in the input sequence and $d_{p}$ is the dimensionality of the phoneme embedding space. Let $\mathbf{E}_{s}\in\mathbb{R}^{T_{t}\times d_{t}}$ denote the sarcasm-aware semantic textual embeddings produced by the fine-tuned LLaMA 3. In the cross-attention module, we treat $\mathbf{E}_{p}$ as the query and $\mathbf{E}_{s}$ as key and value:
@@
where $\mathbf{W}_{q},\mathbf{W}_{k},\mathbf{W}_{v}$ are learned projection matrices, and $d_{k}$ is the dimensionality of the key. The resulting cross-attention output $\mathbf{H}$ represents a phoneme-aligned semantic encoding.
Finally, the prosody exemplars $\mathcal{E}_{w}=\{\mathbf{E}_{w_{1}},\dots,\mathbf{E}_{w_{K}}\}$ extracted by WavLM are linearly projected and added to the cross-attention output over semantic and phoneme embeddings to modulate the decoder:
@@
Experiments
To investigate how semantic and prosodic cues contribute to the perception of sarcasm, we conducted a series of experiments using synthesized speech with systematically manipulated cues. We define a set of experimental conditions that independently and jointly vary semantic and prosodic information. Specifically, we examine:
  • Baseline: speech generated without sarcasm-specific semantic or prosodic guidance;
  • Semantic-only: speech conditioned on sarcasm-aware semantic embeddings derived from a fine-tuned LLaMA 3 model;
  • Prosody-only: speech guided by prosodic exemplars retrieved from a database of sarcastic utterances;
  • Combined: speech incorporating both semantic and prosodic guidance.
In later analyses, we further explore sub-variants within the semantic-only condition, including embeddings from BERT, pretrained LLaMA 3, and LoRA-fine-tuned LLaMA 3, in order to assess the effect of different semantic representations. These conditions collectively allow us to evaluate the relative contributions of semantic cues, prosodic expression, and their interaction to listeners’ perception of sarcastic intent.
Data
For the pre-training phase of VITS, we use the HiFi-TTS corpus [2], a high-fidelity audiobook dataset with diverse speakers and wideband recordings.
@@
To adapt LLaMA 3 as sarcasm-aware semantic encoders via LoRA, we fine-tune them on the News Headlines Sarcasm dataset [22]. This dataset contains over 28,000 headlines from sources such as The Onion, each annotated with a binary sarcasm label. The class distribution is roughly balanced, making it suitable for learning discourse-level markers of sarcastic intent. The LLM serves as a feature extractor that provides high-level semantic embeddings conditioned on sarcastic intent. To validate the effectiveness of this adaptation, we train a sarcasm detection model using various textual embeddings, including BERT, LLaMA 3, LLaMA 3-LoRA, and compare their performance as evidence of each embedding’s ability to capture sarcastic intent.
Experimental Setup
@@
Compared Methods
To systematically assess the contribution of semantic and prosodic cues to sarcasm perception, we synthesized speech under six conditions corresponding to different manipulations of these cues:
  • Baseline: Speech generated using the standard variational text-to-speech model VITS, without any sarcasm-specific semantic or prosodic guidance, serving as a neutral reference.
  • Semantic-only (BERT): Speech conditioned on semantic embeddings from a pretrained BERT model [9], providing general contextual semantic features.
  • Semantic-only (LLaMA 3): Speech conditioned on embeddings from LLaMA 3 [10], enabling richer contextual and pragmatic information.
  • Semantic-only (LLaMA 3-LoRA): Speech conditioned on sarcasm-aware semantic embeddings from a LoRA-fine-tuned LLaMA 3 model, capturing discourse-level cues indicative of sarcastic intent.
  • Prosody-only (RAG): Speech guided by prosodic exemplars retrieved from a database of sarcastic utterances, providing realistic intonation patterns while keeping semantic content neutral.
  • Semantic + Prosody (Proposed): Speech generated using both sarcasm-aware semantic embeddings (LLaMA 3-LoRA) and retrieved prosodic exemplars (RAG), allowing joint manipulation of semantic incongruity and prosodic expression to maximize perceived sarcasm.
Results and Discussion
Effect of Different Semantic Embeddings on Sarcasm Intent Modeling
We first evaluate how well different semantic embeddings capture semantic cues associated with sarcastic intent. Specifically, we build sarcasm detection models and train and evaluate them on the text transcripts of the MUStARD++ dataset, using the same 8:1:1 data split and following the same experimental setup as [25]. The goal of this experiment is to establish a text-level upper bound on sarcasm recognition performance and to motivate the choice of semantic embeddings used later for conditioning speech synthesis. Models are evaluated with Precision (P), Recall (R), and weighted F1-score (F1), with results summarized in Table 1.
Table 1: Performance of different models on sarcasm detection. MUStARD++ (text) refers to the original sarcasm detection systems used in [25].
@@
Both the MUStARD++ baseline and BERT achieved moderate performance with F1-scores around 67%. LLaMA 3 performed slightly worse (65.5%), suggesting that general-purpose pretraining alone is insufficient to capture the nuanced cues of sarcasm. In contrast, PEFT with LoRA yielded a clear improvement, raising the F1-score to 72.5%. This indicates that targeted fine-tuning effectively operationalizes sarcasm-relevant semantic features. These results provide evidence that semantic embeddings from the LoRA-enhanced model more reliably encode discourse-level markers of sarcasm. Importantly, this motivates their use in the TTS framework: by conditioning speech synthesis on embeddings that better reflect sarcastic intent, we can systematically examine whether improvements in semantic cue representation translate to perceptually stronger expressions of sarcasm in synthesized speech.
Sarcastic Speech Synthesis Results
Table 2: Evaluation of sarcastic speech synthesis systems. Objective acoustic metrics (MCD, pitch, energy) assess signal fidelity; sarcasm detection (Precision (P), Recall (R), and weighted F1-score (F1)) evaluates how well sarcastic intent can be recovered from synthesized speech; subjective ratings measure perceived naturalness (NMOS) and sarcasm expressivity (SMOS).
Cue Type Objective Sarcasm Detection Subjective
Method Semantic Prosodic MCD $\downarrow$ Pitch $\downarrow$ Energy $\downarrow$ P $\uparrow$ R $\uparrow$ F1 $\uparrow$ NMOS $\uparrow$ SMOS $\uparrow$
Ground Truth 62.8 62.3 62.3 $3.8\pm 0.1$ $4.5\pm 0.2$
Baseline $9.8$ $261.9$ $4.4$ 60.3 60.5 59.9 $2.6\pm 0.1$ $3.2\pm 0.2$
w/ BERT 9.6 $265.6$ $4.5$ 60.5 60.6 60.5 $2.5\pm 0.1$ $3.1\pm 0.2$
w/ LLaMA 3 $10.4$ $261.6$ $4.4$ 59.6 59.7 59.0 $2.0\pm 0.1$ $2.6\pm 0.2$
w/ LLaMA 3-LoRA $10.1$ $282.6$ $4.4$ 60.6 60.8 60.6 $2.6\pm 0.1$ $3.8\pm 0.2$
w/ RAG $10.0$ $261.0$ $4.4$ 61.5 61.7 61.6 $2.6\pm 0.1$ $3.7\pm 0.2$
w/ LoRA + RAG $9.8$ 259.6 $4.4$ 62.7 62.9 62.5 $\mathbf{2.7\pm 0.1}$ $\mathbf{3.8\pm 0.2}$